OpenStudy (anonymous):

if f(x)=5x-6. what is the equation for f^-1(x)?

OpenStudy (anonymous):

\[y=5x-6\] \[x=5y-6\] \[x+6=5y\] \[\frac{x+6}{5}=y\]

OpenStudy (anonymous):

inverse is \[f^{-1}(x)=\frac{x+6}{5}\]